Consider a hypothetical economy characterized by the following behavioral equations.

IS:                                   LM:

C=400-200r+0.2Y             Md=Y-1000r

I= 240-400r                        Ms=800

G=200                                 P=2

a)derive the equations of the IS and LM curves of the economy

b) Find the equilibrium income level and the interest rate


1
Expert's answer
2021-08-17T16:56:27-0400

(a) Derivation of IS curve equation

Given;

"C=400-200r+0.2Y"

"I=240-400r"

"G=200"


"Y=400-200r+0.2Y+240-400r+200"

"Y-0.2Y=840-600r"

"0.8Y=840-600r"

Thus the IS curve equation is ; "Y=1050-750r"

Derivation of LM curve

Given;

"Md=Y-1000r"

"Ms=800"

"P=2"


"Ms=Mt+Mz"

"800=Y-1000r+2"

"800+1000r-2=Y"

"798+1000r=Y"

The LM curve equation is ; "Y=798+1000r"


(b)

To find interest rate , equate IS curve equation with LM curve equation

"1050-750r=798+1000r"

"1050-798=1000r+750r"

"r=0.144= 14.4%" %

To find equilibrium income, substitute the value of r in the IS curve equation:

"Y=1050-750r"

"Y=1050-750(0.144)"

"Y=942"